OpenStudy (cicilybailey14):

Fan+Medal! Suppose y varies inversely with x. Write an equation for the inverse variation. y = 6 when x = 1/3 xy = 1/18 xy = 1/3 xy = 2 xy = 1/2

OpenStudy (michele_laino):

please note that: x*y= \[=6*\frac{ 1 }{ 3 }=2=k\] so x*y=2